OpenStudy (anonymous):

An unknown number y is 15 more than an unknown number x. The number y is also x less than 8. The equations to find x and y are shown below. y = x + 15 y = −x + 8 Which of the following statements is a correct step to find x and y? Multiply the equations to eliminate y. Add the equations to eliminate x. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

OpenStudy (anonymous):

Which one do you think it is?

OpenStudy (anonymous):

Multiply the equations to eliminate y.

OpenStudy (anonymous):

Well if you multiply the equations, y won't be eliminated, so that can't be it.

OpenStudy (anonymous):

Write the points where the graphs of the equations intersect the y-axis.??

OpenStudy (anonymous):

No. You would actually want where the two lines intersect.

OpenStudy (hybrik):

"Add the equations to eliminate x." is correct

OpenStudy (hybrik):

To solve this you need to combine the equation into one

OpenStudy (hybrik):

So you add X to remove X to find Y

OpenStudy (hybrik):

Which gives you y = 23, then substitute Y into any equation to find X

OpenStudy (anonymous):

@hybrik I was getting to that. On Openstudy, you want to help them find the answer them-self, not give them the answer.

OpenStudy (anonymous):

Betsey is 11 years older than Waldo. Betsey's age is 15 years less than three times Waldo's age. The system below models the relationship between Betsey's age (b) and Waldo's age (w): b = w + 11 b = 3w − 15 Which of the following methods is correct to find Betsey's and Waldo's age? Solve w + 11 = 3w − 15 to find the value of w. Solve b + 11 = 3b − 15 to find the value of b. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

OpenStudy (anonymous):

Again, which do you think it is?

OpenStudy (anonymous):

Solve w + 11 = 3w − 15 to find the value of w.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Bella and Heather put some money into their money boxes every week. The amount of money (y), in dollars, in their money boxes after a certain amount of time (x), in weeks, is shown by the equations below: Bella: y = 25x + 60 Heather: y = 30x + 10 After how many weeks will Bella and Heather have the same amount of money in their money boxes, and what is that amount? 10 weeks, $10 10 weeks, $310 9 weeks, $310 310 weeks, $10

OpenStudy (anonymous):

I think its 10,10

OpenStudy (hybrik):

@ComputerNerd I just started 20 minutes ago, sorry for not knowing..

OpenStudy (anonymous):

I need helpppppppppppppppppppppp

OpenStudy (anonymous):

@Chumpian

OpenStudy (anonymous):

Variable x is 7 more than variable y. Variable x is also 1 less than y. Which of the following pairs of equations best models the relationship between x and y? x = 1y x = y + 7 x = y + 7 x = y − 1 y = x + 8 y = x − 1 y = 7x y = x + 1